the length of a rectangle is 5 more than the width. the area is 336 square meters. find the length and the width of the rectangle

Answer by checkley79(3341)

AREA=LW
L=W+5
336=(W+5)W
336=W^2+5W
W^2+5W-336=0
(W+21)(W-16)=0
W-16=0
W=16 ANS. FOR THE WIDTH.
L=16+5=21 ANS. FOR THE LENGTH.
PROOF:
336=21*16
336=336